# Transforming Code Repositories into MCP Services

### v2

@@ 1 Introduction
The landscape of artificial intelligence is increasingly defined by autonomous agents that leverage Large Language Models (LLMs) to interact with external tools (Wang et al., 2024; Xi et al., 2024; Bubeck et al., 2023). To overcome the inherent limitations of LLMs in tasks requiring real-time information or precise computation, the paradigm of tool-augmented reasoning has become central (Huang et al., 2024; Hao et al., 2023; Yue et al., 2024). Seminal works have demonstrated that models can effectively learn to invoke external functions (Schick et al., 2023; Qin et al., 2023; Parisi et al., 2022).
@@
In response to this integration challenge, the community’s efforts have evolved, inadvertently revealing a deeper, more foundational bottleneck (Yue et al., 2025a). The initial challenge is to establish the fundamental feasibility of tool use, where a limited set of tools proves the feasibility of the paradigm (Schick et al., 2023; Patil et al., 2023; Ding et al., 2025). To break past the inherent scarcity of these platforms, the focus shifts to the vast landscape of open-source repositories (Wang et al., 2025; Xie et al., 2023). This move, however, trades a scarcity problem for a chaos problem, exposing the wild non-standardization of real-world code. The MCP emerges as a direct answer to this chaos, promising a universal interface (Zhang et al., 2025; Anthropic, 2023). Yet, this leads to the critical gap: research is focused almost exclusively on the consumption side of MCP, using services from a presumed-to-exist pool (Gan & Sun, 2025), while the foundational “supply-side” problem of how to populate this pool from existing software is largely unaddressed.
@@
However, while these efforts advance the consumption side of the problem, how agents can better use tools, they largely overlook a more fundamental bottleneck on the supply side. This supply bottleneck is not a theoretical concern but a stark reality preventing standards like MCP from achieving widespread adoption. For example, RAG-MCP (Gan & Sun, 2025) utilizes over 4,400 servers on mcp.so, but there are 268 million public GitHub repositories. The critical question of how to create a large and diverse pool of these standardized, agent-ready tools has been left unaddressed. This creates a major adoption gap, effectively locking away the largest software repository, GitHub, from this emerging ecosystem.
In this paper, we introduce Code2MCP, a new framework designed to bridge the critical tool supply gap. Code2MCP presents a blueprint for transforming any GitHub repository into a functional and documented MCP service with minimal human intervention. However, this transformation is a complex endeavor encompassing four pivotal challenges: (1) deep code comprehension to identify core functionalities, (2) reliable environment replication to ensure executability, (3) intelligent tool abstraction to design useful and valid service interfaces, and (4) robust self-correction to handle the inevitable errors throughout the process. To systematically address these challenges, as shown in Figure 2, Code2MCP implements a collaborative multi-agent system (Park et al., 2023). Unlike general-purpose coding agents (Cognition, 2024; Jimenez et al., 2024), different agents in our framework are specialized for the distinct stages of code analysis, environment setup, and API design. Crucially, the overall reliability of this workflow is ensured by an integrated Run-Review-Fix self-correction cycle, which endows the system with the ability to autonomously debug and refine the entire conversion process. The key contributions of this work are listed as follows:
  • To solve the fundamental tool supply bottleneck hindering the adoption of the MCP standard, we propose a novel automated framework, Code2MCP, the first framework to systematically transform code repositories into agent-ready MCP services.
@@
Consumer
Toolformer Teaching LLM to use external tools 5 predefined tools Fine-tuning $\times$
SciToolAgent
Domain-specific enhancement
for scientific tool utilization
KG of Scientific Tools
(500+ tools)
Retrieval on KG $\times$
HuggingGPT
Increase the size
of tool pool
huggingface.co LLM task planning $\times$
Gorilla
TorchHub, TensorHub
(1600+ tools)
Retriever-aware
training
$\times$
OpenAgents
Tool use from open-source
beyond a closed pool
github.com Multi-agent planning $\times$
@@ 2 Related Work
Exploring Unstructured Open-Source Repositories and Challenges. A more ambitious paradigm shift involves treating the entirety of open-source code repositories as a virtually infinite tool source. Frameworks like OpenAgents (Xie et al., 2023) and RepoMaster (Wang et al., 2025) empower agents to directly parse, reason about, and execute code within GitHub repositories. These works confront the complexity of real-world code but expose the core bottleneck: the vast majority of these repositories are not designed for programmatic use by LLM agents. They lack standardized interfaces (Zhang et al., 2024; Jin et al., 2024; Ray, 2025), forcing the agent into an ad-hoc, brittle, and unreliable process of reverse-engineering the code, setting up its environment, and managing dependencies for every single task (Zeng et al., 2024; Olausson et al., 2023). This chaotic integration process illustrates a critical failure on the tool supply side.
The Emergence of Standardization and Unaddressed Gap. Recognizing this chaos, the community has moved towards standardization, exemplified by the MCP. For example, RAG-MCP (Gan & Sun, 2025) explores how an agent can effectively retrieve and select the most appropriate MCP service from a pool of 4400+ available options. This approach is promising, but it presumes the existence of a rich ecosystem of MCP-compliant services (Hasan et al., 2025). This highlights a crucial, unaddressed gap: how is this ecosystem of MCP services populated in the first place?
@@ 3 Methodology: The Code2MCP Framework
To achieve the goal of automatically transforming an arbitrary GitHub repository into a fully functional and reliable MCP service, we design Code2MCP, an automated framework driven by the collaboration of seven specialized agents. The entire conversion process, as depicted in Figure 2, is a multi-agent workflow that begins with code analysis, proceeds through a core Run-Review-Fix self-correction loop, and culminates in the generation of a merge-ready pull request.

### v3

@@ 1 Introduction
The landscape of artificial intelligence is increasingly defined by autonomous agents that leverage Large Language Models (LLMs) to interact with external tools (Wang et al., 2024; bubeck2023sparks). To overcome the inherent limitations of LLMs in tasks requiring real-time information or precise computation, the paradigm of tool-augmented reasoning has become central (Huang et al., 2024; Hao et al., 2023). Seminal works have demonstrated that models can effectively learn to invoke external functions (Schick et al., 2023; Qin et al., 2023; Parisi et al., 2022).
@@
In response to this integration challenge, the community’s efforts have evolved, inadvertently revealing a deeper, more foundational bottleneck (Yue et al., 2025). The initial challenge is to establish the fundamental feasibility of tool use, where a limited set of tools proves the feasibility of the paradigm (Schick et al., 2023; Patil et al., 2023; Ding et al., 2025; Qin et al., 2024). To break past the inherent scarcity of these platforms, the focus shifts to the vast landscape of open-source repositories (Wang and others, 2025; Xie et al., 2023; Wang, 2023). This move, however, trades a scarcity problem for a chaos problem, exposing the wild non-standardization of real-world code. The MCP emerges as a direct answer to this chaos, promising a universal interface (Yu et al., 2025; Anthropic, 2023). Yet, this leads to the critical gap: research is focused almost exclusively on the consumption side of MCP, using services from a presumed-to-exist pool (Gan and Sun, 2025), while the foundational “supply-side” problem of how to populate this pool from existing software is largely unaddressed.
@@
However, while these efforts advance the consumption side of the problem, how agents can better use tools, they largely overlook a more fundamental bottleneck on the supply side. This supply bottleneck is not a theoretical concern but a stark reality preventing standards like MCP from achieving widespread adoption. For example, RAG-MCP (Gan and Sun, 2025) utilizes over 4,400 servers on mcp.so, but there are 268 million public GitHub repositories. The critical question of how to create a large and diverse pool of these standardized, agent-ready tools has been left unaddressed. This creates a major adoption gap, effectively locking away the largest software repository, GitHub, from this emerging ecosystem.
In this paper, we introduce Code2MCP, a new framework designed to bridge the critical tool supply gap. Code2MCP presents a blueprint for transforming any GitHub repository into a functional and documented MCP service with minimal human intervention. However, this transformation is a complex endeavor encompassing four pivotal challenges: (1) deep code comprehension to identify core functionalities, (2) reliable environment replication to ensure executability, (3) intelligent tool abstraction to design useful and valid service interfaces, and (4) robust self-correction to handle the inevitable errors throughout the process. To systematically address these challenges, as shown in Figure 2, Code2MCP implements a collaborative multi-agent system (Park et al., 2023). Unlike general-purpose coding agents (Cognition, 2024; Jimenez et al., 2024; Zhang et al., 2024b; Huang et al., 2023), different agents in our framework are specialized for the distinct stages of code analysis, environment setup, and API design. Crucially, the overall reliability of this workflow is ensured by an integrated Run-Review-Fix self-correction cycle, which endows the system with the ability to autonomously debug and refine the entire conversion process. The key contributions of this work are listed as follows:
  • To solve the fundamental tool supply bottleneck hindering the adoption of the MCP standard, we propose Code2MCP, a novel automated framework that, to the best of our knowledge, is the first to systematically transform code repositories into agent-ready MCP services.
@@
Consumer Toolformer Teaching LLM to use external tools 5 predefined tools Fine-tuning $\times$
SciToolAgent Domain-specific enhancement for scientific tool utilization KG of Scientific Tools (500+ tools) Retrieval on KG $\times$
HuggingGPT Increase the size of tool pool huggingface.co LLM task planning $\times$
Gorilla TorchHub, TensorHub (1600+ tools) Retriever-aware training $\times$
OpenAgents Tool use from open-source beyond a closed pool github.com Multi-agent planning $\times$
@@ 2 Related Work
Exploring Unstructured Open-Source Repositories and Challenges. A more ambitious paradigm shift involves treating the entirety of open-source code repositories as a virtually infinite tool source. Frameworks like OpenAgents (Xie et al., 2023) and RepoMaster (Wang and others, 2025) empower agents to directly parse, reason about, and execute code within GitHub repositories. These works confront the complexity of real-world code but expose the core bottleneck: the vast majority of these repositories are not designed for programmatic use by LLM agents. They lack standardized interfaces (Zhang et al., 2024a; Jin et al., 2024; Ray, 2025), forcing the agent into an ad-hoc, brittle, and unreliable process of reverse-engineering the code, setting up its environment, and managing dependencies for every single task. This chaotic integration process illustrates a critical failure on the tool supply side.
The Emergence of Standardization and Unaddressed Gap. Recognizing this chaos, the community has moved towards standardization, exemplified by the MCP. For example, RAG-MCP (Gan and Sun, 2025) explores how an agent can effectively retrieve and select the most appropriate MCP service from a pool of 4400+ available options. This approach is promising, but it presumes the existence of a rich ecosystem of MCP-compliant services (Hasan et al., 2025). This highlights a crucial, unaddressed gap: how is this ecosystem of MCP services populated in the first place?
@@ 3 Methodology: The Code2MCP Framework
To achieve the goal of automatically transforming an arbitrary GitHub repository into a fully functional and reliable MCP service, we design Code2MCP, an automated framework driven by the collaboration of seven specialized agents. The entire conversion process, as depicted in Figure 2, is a multi-agent workflow that begins with code analysis, proceeds through a core Run-Review-Fix self-correction loop, and culminates in the generation of a merge-ready pull request. For completeness, a more formal description of this workflow is provided in the appendix A.3.
@@
Initialization and Analysis. The Download Agent first clones the specified repository, identified by its URL $u$, into an isolated local workspace. The Environment Agent then replicates the runtime environment from dependency files or Dockerfiles, addressing one of the most common failure points in code conversion and supporting reliable subsequent code generation and testing.
Once the environment is ready, the Analysis Agent identifies tool-worthy functionalities within the codebase. It leverages the DeepWiki tool to obtain a semantic view of the code and associates code entities with their intent from documentation and comments. The output is a Code Report that summarizes candidate APIs and guides subsequent stages.
Generation, Execution, and Self-Correction. Given this conversion blueprint, the framework enters its core iterative loop that turns the identified functionalities into executable MCP services.
The loop starts with the Generation Agent, which takes the Code Report and uses an LLM to abstract the core functionalities into MCP-compliant interfaces. It creates the tool interface definitions and adapter file that connect the original code to the MCP interface, together with a basic test suite.
Once the code is generated, the Run Agent executes the test suite in the prepared environment to verify executability. If the tests pass, the workflow proceeds to finalization; otherwise, the Run Agent records the error traceback $\tau$ and forwards it to the Review Agent. The Review Agent analyzes $\tau$ together with the generated code, the Code Report, and the failing test, and diagnoses root causes such as logic errors, missing dependencies, or interface mismatches. It then formulates a correction plan $\delta$ that specifies which files and code blocks to change, and hands this plan back to the Generation Agent to re-synthesize the MCP files. This Run-Review-Fix loop repeats until tests pass or a maximum of $B$ attempts is reached.
Finalization and Delivery. After the core loop succeeds, the Finalize Agent organizes and packages the validated MCP service files. To facilitate review and adoption by the original repository maintainers, it also generates a README file explaining how to use the new MCP service. All artifacts are arranged into a reproducible directory structure under the workspace (Appendix A.2), and the agent can prepare a pull request to submit these additions back to the original repository.
4 Evaluation
4.1 Experimental Setup
Task. The evaluation task is: given a GitHub repository URL, Code2MCP automatically generates an MCP service for that repository, and the outcome of the conversion is judged as a success or a failure. The evaluation focuses on the overall conversion success rate across different domains and repository characteristics, and on the time and stability of Code2MCP compared with manual implementation and a GPT-4 template-based baseline on the same set of repositories.
Repositories. Code2MCP is evaluated on 50 open-source repositories from 10 scientific and engineering domains (5 per domain), covering biomedical science, psychology, mathematical computing, earth science, chemistry, physics, astronomy, social science, linguistics, and econometrics. These repositories range from small utility libraries to large frameworks with complex dependency stacks. For each repository, the evaluation logs the environment construction result, MCP conversion result, number of generated tools, and failure types observed across all Run-Review-Fix rounds; per-repository details and the associated failure labels are summarized in a large table in the Appendix C, and a subset of representative scientific-computing repositories is further used for the Run-Review-Fix ablation and qualitative case studies.
Success criterion. A conversion is regarded as successful only if the following three conditions all hold: the runtime environment can be reconstructed according to the dependency specifications provided by the repository, the generated MCP server can be started and must pass a unified RunNode test, and at least three core tools can be invoked correctly in automatic tests and return valid outputs. If any of these conditions is not satisfied, the repository is counted as a failure.
Failure taxonomy. To analyze failure modes, all failed repositories are annotated with one or more labels from the following six categories:
  • env_failure. The environment or dependencies cannot be reconstructed.
  • api_inference_error. Systematic errors occur when inferring tool interfaces.
  • import_error. Import paths or cross-module dependencies are misconfigured.
  • repo_internal_bug. Bugs or version conflicts inside the original repository prevent it from running reliably even in its native environment.
  • mcp_spec_violation. The generated JSON Schema or response structure does not conform to the MCP specification.
  • untoolable_repo. The repository is a collection of scripts or heavily depends on interactive CLIs or GUIs, which are difficult to abstract into stable MCP tools.
Table 2: Per-domain summary of environment setup success, basic test success, recovery by the Run-Review-Fix (RRF) loop (“–” if none are recovered), and final MCP conversion success.
Domain Repos Env succ. Test succ. RRF recovered Avg rounds MCP succ.
Biomedical 5 4/5 (80%) 2/5 (40%) 1/2 (50%) 1.5 3/5 (60%)
Psychology 5 3/5 (60%) 3/5 (60%) 0/1 (0%) 3/5 (60%)
Math 5 5/5 (100%) 3/5 (60%) 1/2 (50%) 2.0 4/5 (80%)
Earth Science 5 3/5 (60%) 2/5 (40%) 1/3 (33.3%) 1.0 3/5 (60%)
Chemistry 5 4/5 (80%) 2/5 (40%) 1/2 (50%) 1.3 3/5 (60%)
Physics 5 3/5 (60%) 2/5 (40%) 0/2 (0%) 2/5 (40%)
Astronomy 5 4/5 (80%) 3/5 (60%) 1/1 (100%) 1.5 4/5 (80%)
Social Science 5 4/5 (80%) 2/5 (40%) 1/2 (50%) 1.0 3/5 (60%)
Linguistics 5 5/5 (100%) 3/5 (60%) 1/2 (50%) 2.0 4/5 (80%)
Econometrics 5 3/5 (60%) 2/5 (40%) 1/2 (50%) 1.0 3/5 (60%)
Overall 50 38/50 (76%) 24/50 (48%) 8/19 (42.1%) 1.4 32/50 (64%)
4.2 Large-scale Repository Conversion and Failure Analysis
Overall and per-domain success. The overall conversion success rate is first analyzed at the domain level. In Table 2, for each domain, the number of repositories with successful environment setup and basic test passing, the number of initially failing repositories recovered by the Run-Review-Fix loop and their average rounds, and the final number of successful MCP conversions. The per-domain MCP success rates range from 0.40 to 0.80: domains dominated by library-style projects, such as mathematical computing, astronomy, and linguistics, tend to have higher success rates, while domains such as physics and econometrics, which contain more complex environments and mixtures of scripts and workflow-style code, tend to have lower success rates. All statistics in the table are computed on a per-repository basis, where each repository contributes a single final outcome after the Run-Review-Fix loop. Per-repository results, including environment construction and MCP conversion outcomes as well as the number of generated tools, are summarized in the large table in the appendix and can be further grouped by domain or repository size if desired.
Figure 3: Distribution of 42 failure labels assigned to the 18 failed repositories. A single repository can trigger multiple failure types across different Run-Review-Fix rounds.
Failure modes. To understand why conversions fail on real repositories, all 18 failed repositories in the 50-repository evaluation are annotated using the six failure labels introduced in the failure taxonomy. Labels are aggregated across all Run-Review-Fix rounds, so a single repository may receive multiple labels if different error types are observed during different stages of the pipeline. The distribution of these labels is visualized in Figure 3. The labels env_failure and api_inference_error account for more than half of all failures, indicating that environment reconstruction and interface inference are the primary bottlenecks of the current workflow. The labels import_error and untoolable_repo are also relatively common, reflecting the difficulty of handling complex import paths, initialization order, and script-driven repositories in a fully automated pipeline.
Run-Review-Fix ablation. To assess the specific contribution of the Run-Review-Fix self-correction loop, an ablation study is conducted on a set of representative scientific-computing repositories by comparing two configurations: a single-pass configuration without the loop and the full Code2MCP pipeline with the loop enabled. Figure 4 summarizes the results. The bar chart on the left shows that the per-repository success rates increase for most repositories once the loop is enabled, often moving from medium success under a single-generation setting to success rates close to or on par with human-written wrappers. The line chart on the right plots the average number of remaining errors as a function of the number of Run-Review-Fix rounds, where “remaining errors” denote the number of failing assertions or uncaught exceptions in the unified test after each round. These results indicate that the majority of initial failures can be automatically repaired within one to three Run-Review-Fix rounds, whereas the few repositories that still fail after multiple rounds represent the main current limitation of the method.
4.3 Comparison with Human and GPT-4 Baselines
The evaluation compares three configurations on the same set of representative repositories:
  • a human configuration, where MCP wrappers are implemented from scratch by human developers;
  • a GPT-4 template configuration, where a single GPT-4 agent generates MCP wrappers;
  • the Code2MCP configuration, which uses the multi-agent pipeline with the Run-Review-Fix loop.
Figure 4: Repo success rates with and without Run‑Review‑Fix (left) and remaining errors across RRF rounds (right) for five representative scientific MCP repositories.
In the human configuration, ten graduate students with at least three years of Python experience and basic familiarity with MCP concepts implement wrappers from scratch. Before the study, all participants read the MCP specification and a commented example repository. Each participant is randomly assigned several repositories, and the measured time is from the moment they start reading the repository to the point where the minimal test passes.
In the GPT-4 template configuration, a unified system prompt instructs a single GPT-4 model to generate a complete MCP service implementation based on the repository README and several key source files selected by static analysis. We use the same base model and decoding configuration as in Code2MCP, namely gpt-4o-2024-05-13 with temperature set to 1.0, to ensure a fair comparison. The model is allowed up to three dialogue turns: the first turn generates an initial implementation, and up to two additional turns may revise the code based on execution errors returned by a test runner. No additional agents or planning mechanisms are used, and the full system prompt is provided in Appendix B.1.
Figure 5: Average task success rate (left) and average completion time (right) across the ten scientific domains for the three configurations: Human experts, GPT-4 with basic scaffolding, and Code2MCP.
In the Code2MCP configuration, the default multi-agent pipeline with the Run-Review-Fix loop is used. For each repository, we record the final success status and the completion time, defined as the end-to-end wall-clock time starting from the beginning of environment setup and ending when the unified MCP test first passes or the run is terminated; see Appendix B.2 for details. Figure 5 summarizes the average task success rates and completion times of the three configurations across the ten scientific domains. For each configuration and domain, task success rate is defined as the fraction of successful attempts over all attempts on representative repositories in that domain. Completion time is the average end-to-end wall-clock time per attempt, measured from the beginning of environment setup until the minimal MCP test first passes or the run is terminated.
Figure 6: Task success rates of Human experts, GPT-4 with a template, and Code2MCP on representative repositories in the four repository groups. The x-axis lists representative repositories in each group, labeled by domain abbreviation and index (e.g., Math-1).
The overall trend is as follows: Code2MCP achieves task success rates close to human implementations in most domains, and consistently higher than the GPT-4 template configuration. In terms of completion time, human implementations are typically on the order of hours, while Code2MCP completes in tens of minutes; the GPT-4 template configuration is the fastest but has lower success rates.
When repositories are grouped by dependency complexity and project type into low-dependency libraries, medium-dependency scientific libraries, high-dependency libraries, and script-heavy projects, Figures 6 and 7 further break down the same comparison on representative repositories within each group. The bucketing rules are detailed in Appendix B.2.
Figure 7: Completion time of Human experts, GPT-4 with a template, and Code2MCP on representative repositories in the four repository groups.
4.4 Integration with Existing Tool Systems
This subsection summarizes how Code2MCP integrates with existing tool ecosystems by combining it with RepoMaster and OpenAgents on a shared set of tool-usage tasks.
RepoMaster + Code2MCP. In the RepoMaster setting, code-understanding and tool-usage tasks are run in two configurations: using RepoMaster alone to operate at the source-code level, and using RepoMaster together with Code2MCP, where RepoMaster first selects relevant repositories and then invokes the corresponding MCP tools. Figure 8 shows that the configuration with Code2MCP achieves higher task success rates and requires fewer interaction steps (see Appendix B.2 for metric definitions), suggesting that delegating part of the work to standardized MCP tools improves efficiency without modifying the upstream agent policy.
OpenAgents + Code2MCP. In the OpenAgents setting, cross-domain tasks are evaluated under two configurations: using OpenAgents alone, and using OpenAgents with MCP tools automatically generated by Code2MCP added to the tool pool. As shown in Figure 9, we report the task success rate and coverage, where coverage and success rate are defined in Appendix B.2. Adding Code2MCP tools increases both metrics across all task groups, indicating that a richer MCP tool pool can be effectively exploited by existing planning and retrieval strategies.
Overall, RepoMaster and OpenAgents handle tool discovery and code retrieval on the consumer side, while Code2MCP supplies additional MCP services on the supply side. Under the same front-end selection and planning strategies, integrating Code2MCP enriches the tool pool and yields measurable improvements in end-to-end task performance.
Figure 8: RepoMaster with and without Code2MCP tools: task success rate (left) and average interaction steps (right) across task groups.
4.5 Case Studies: Protein, Math, and Computational Fluid Dynamics
This subsection highlights three representative repositories from the 50-repository evaluation set: the biomedical protein modeling library ESM, the symbolic mathematics library SymPy, and the CFD framework Foam-Agent built on OpenFOAM, illustrating the kinds of MCP tools Code2MCP generates and how agents use them in practice.
Figure 9: OpenAgents with and without Code2MCP tools: task success rate (left) and coverage (right) across general and scientific task groups.
Bridging protein science with AI agents. In protein science, models such as AlphaFold (Jumper et al., 2021) greatly improve structure prediction, but using them in everyday research typically requires substantial scripting and environment setup. ESM models complement AlphaFold with efficient sequence modeling and zero-shot variant effect prediction, yet traditionally still require multi-step Python pipelines for computing physicochemical properties, predicting structures, and analyzing mutations. Code2MCP converts ESM into an MCP service by exposing tools such as Analyze_sequence, Predict_structure, and Predict_variant_effect. We provide a qualitative example of using ESM-based MCP tools for protein analysis in Appendix A.5.
Figure 10: An AI agent orchestrates a CFD simulation pipeline by invoking Generate_mesh, Generate_simulate, and Visualize_velocity functions to guide a user through the entire process, from initial mesh generation and simulation setup to the final velocity visualization.
Enhancing mathematical reliability in AI agents. For mathematical computing, large language models are prone to errors in symbolic derivations and exact calculations, whereas libraries such as SymPy already provide reliable implementations. Code2MCP organizes common capabilities from SymPy into a mathematical MCP service with tools for limits and integration, matrix operations, symbolic simplification, and transforms such as Fourier and Laplace. When asked to compute an integral or the volume of a solid of revolution, an agent simply invokes the corresponding tool and returns an exact answer, relying on SymPy as a backend instead of ad-hoc prompt-based reasoning.
Automating CFD simulation for AI agents. CFD workflows typically involve mesh generation, solver configuration, execution, and post-processing, traditionally coordinated by engineers through shell scripts and configuration files. Foam-Agent provides a higher-level interface to OpenFOAM, and Code2MCP further turns it into an MCP service with tools such as Generate_mesh, Generate_simulate, and Visualize_velocity. As illustrated in Figure 10, a user describes the target geometry and physical parameters in natural language, and the agent calls these tools in sequence to construct the mesh, run the simulation, and visualize the velocity field.
5 Conclusion
Addressing the key challenge of insufficient tool supply in the MCP ecosystem, this paper introduces Code2MCP, a framework that automatically converts GitHub repositories into functional MCP services. The framework employs a multi-agent workflow for code comprehension, environment reconstruction, and tool abstraction, augmented by a “Run-Review-Fix” self-correction loop that improves end-to-end reliability across diverse scientific libraries. Our evaluation on 50 repositories shows that Code2MCP can robustly expose high-value APIs as MCP tools while also revealing clear failure modes and boundary conditions. This work presents an automated pathway for connecting existing codebases to the agent tool ecosystem, and we leave systematically hardening security and building standardized large-scale MCP benchmarks from these converted services as important directions for future work.
Appendix A Appendix
A.1 Use of Large Language Models
A.2 Output Directory Structure
Figure 11 shows the final directory layout generated by Code2MCP, where the original repository and the synthesized MCP artifacts are organized under a unified workspace.
A.3 Detailed Conversion Pipeline and Algorithm
This section outlines the roles of the specialized agents within the Code2MCP framework and their coordination as shown in Algorithm 1, where Download and Environment initialize the workspace, Analysis and Generation construct the MCP service, Run and Review form the Run-Review-Fix loop, and Finalize produces the final deliverables. The following are placeholders for their system prompts.